OpenStudy (anonymous):

Solve the following equation 20 - 3x = 8 x=

OpenStudy (anonymous):

x=4 (:

OpenStudy (31356):

20-3x=8 -20 -20 3x=-12 3x 3x X=4

OpenStudy (31356):

Medal? Sorry I typed too late.

OpenStudy (31356):

Thanks for the medal!